# Scalar tidal response of a rotating BTZ black hole

### v1

@@ 6 Discussion
In this work, we studied the tidal response of a rotating (non-extremal and extremal) BTZ black hole to the scalar tidal perturbation. We found that the scalar tidal response function of a non-extremal rotating BTZ black hole is non-zero and has a real part, consequently implying non-zero tidal Love numbers. However, the tidal response function also vanishes in some cases (listed in 3). Furthermore, static tidal Love numbers of a non-extremal rotating BTZ black hole to a scalar tidal field are generally non-vanishing. However, they vanish for an axi-symmetric tidal perturbation.
Since the transformed coordinate defined for the calculation of the scalar tidal response function of a non-extremal rotating BTZ black hole (7) diverges for an extremal rotating BTZ black hole $(r_{+}=r_{-})$, the calculation presented in 3 is not valid for it. Therefore, we separately calculated the scalar tidal response function for an extremal rotating BTZ black hole with a redefined coordinate transformation equation (21). Qualitatively, the tidal response function of an extremal rotating BTZ black hole shows similar behaviour to that of a non-extremal rotating BTZ black hole. It is non-zero and has a real part, which implies non-zero tidal Love numbers. Here too, the tidal response function vanishes for certain cases (listed in 4). In addition, static tidal Love numbers of an extremal rotating BTZ black hole to the scalar tidal perturbation are generally non-zero. However, they vanish for an axi-symmetric tidal perturbation too.
@@
Finally, we also proposed a procedure to calculate the tidal response function for a charged rotating BTZ black hole. Due to the logarithmic term in the metric of the charged BTZ black hole, we couldn’t calculate the tidal response function in a similar way as done in 3 and 4. Therefore, we performed our calculation in the near-horizon regime, which is an assumption for this specific calculation and has been used in other studies of the tidal response function [11, 24, 27]. With this assumption, the results for a charged rotating BTZ black hole will be similar to the results found in 3 and 4 with some replacements defined in 5 ($(\{r_{\pm},\,M,\,l\}\to\{\tilde{r}_{\pm},\,\tilde{M},\,\tilde{l}\})$. The suggested procedure is also applicable for a charged, non-rotating BTZ black hole, if we set $J=0$.
@@ Acknowledgements
We would like to thank Sumanta Chakraborty for suggesting this problem, and for his help throughout the project.
@@ A.1 Transforming the radial Klein-Gordon equation
where $f(r)$ and $\Omega(r)$ are defined in 2. If we apply the coordinate transformation (7), which is $z=(r^{2}-r_{+}^{2})/(r_{+}^{2}-r_{-}^{2})$, then 39 transforms to
@@
therefore, the coefficient of $R^{\prime}(z)$ can be reexpressed as:
@@
Therefore, 40 becomes (using 42 and 45)
@@ A.2 Solution of the transformed radial Klein-Gordon equation
by choosing the suitable values of $\alpha$ and $\beta$. The above substitution transforms 49 to
@@
If we take any values of $\alpha$ and $\beta$ which satisfy 52 (say $[\alpha_{0},\beta_{0}]$), then 51 becomes
@@ B.1 Transforming the radial Klein-Gordon equation
By using $r^{2}=r_{+}^{2}(1+\tilde{z})$ and the form of $f(r)$ and $\Omega(r)$ (from 61), the coefficient of $R(\tilde{z})$ becomes
@@ B.2 Solution of the transformed radial Klein-Gordon equation
If we take any values of $\alpha$ and $\beta$ which satisfy 70 say $[\alpha_{0},\beta_{0}]$, then the 69 becomes

### v2

@@ 6 Discussion
In this work, we studied the tidal response of a rotating (non-extremal and extremal) BTZ black hole to the scalar tidal perturbation. We found that the scalar tidal response function of a non-extremal rotating BTZ black hole is non-zero and has a real part, consequently implying non-zero TLNs. However, the tidal response function also vanishes in some cases (listed in 3). Furthermore, static TLNs of a non-extremal rotating BTZ black hole to a scalar tidal field are generally non-vanishing. However, they vanish for an axi-symmetric tidal perturbation.
Since the transformed coordinate defined for the calculation of the scalar tidal response function of a non-extremal rotating BTZ black hole (7) diverges for an extremal rotating BTZ black hole $(r_{+}=r_{-})$, the calculation presented in 3 is not valid for it. Therefore, we separately calculated the scalar tidal response function for an extremal rotating BTZ black hole with a redefined coordinate transformation equation (21). Qualitatively, the tidal response function of an extremal rotating BTZ black hole shows similar behaviour to that of a non-extremal rotating BTZ black hole. It is non-zero and has a real part, which implies non-zero TLNs. Here too, the tidal response function vanishes for certain cases (listed in 4). In addition, static TLNs of an extremal rotating BTZ black hole to the scalar tidal perturbation are generally non-zero. However, they vanish for an axi-symmetric tidal perturbation too.
@@
Finally, we also proposed a procedure to calculate the tidal response function for a charged rotating BTZ black hole. Due to the logarithmic term in the metric of the charged BTZ black hole, we couldn’t calculate the tidal response function in a similar way as done in 3 and 4. Therefore, we did the calculation for a slowly varying tidal field in the near-zone region for a BTZ black hole with a small charge. With this assumption, the results for a charged rotating BTZ black hole will be similar to the results found in 3 and 4 with some replacements defined in 5 ($(\{r_{\pm},\,M,\,l\}\to\{\tilde{r}_{\pm},\,\tilde{M},\,\tilde{l}\})$. The suggested procedure is also applicable for a charged, non-rotating BTZ black hole, if we set $J=0$.
@@ Acknowledgements
We would like to thank Sumanta Chakraborty for suggesting this problem, and for his help throughout the project. R. P. B. would also like to thank Sukanta Bose for his support.
@@ A.1 Transforming the radial Klein-Gordon equation
where $f(r)$ and $\Omega(r)$ are defined in 2. If we apply the coordinate transformation (7), which is $z=(r^{2}-r_{+}^{2})/(r_{+}^{2}-r_{-}^{2})$, then 41 transforms to
@@
the coefficient of $R^{\prime}(z)$ can be reexpressed as:
@@
Therefore, 42 becomes (using 44 and 47)
@@ A.2 Solution of the transformed radial Klein-Gordon equation
by choosing the suitable values of $\alpha$ and $\beta$. The above substitution transforms 51 to
@@
If we take any values of $\alpha$ and $\beta$ which satisfy 54 (say $[\alpha_{0},\beta_{0}]$), then 53 becomes
@@ B.1 Transforming the radial Klein-Gordon equation
Using $r^{2}=r_{+}^{2}(1+\tilde{z})$ and the form of $f(r)$ and $\Omega(r)$ (from 63), the coefficient of $R(\tilde{z})$ becomes
@@ B.2 Solution of the transformed radial Klein-Gordon equation
If we take any values of $\alpha$ and $\beta$ which satisfy 72 say $[\alpha_{0},\beta_{0}]$, then the 71 becomes
@@
Appendix C Calculation for charged BTZ black hole
In this appendix, we will chalk out the steps involved in arriving at the approximate form of $f(r)$ and $\Omega(r)$ in 37 for a charged rotating BTZ black hole.
For reference, we are rewriting the form of $f(r)$ and $\Omega(r)$ for charged rotating BTZ black hole (32):
$f(r)=-M+\frac{r^{2}}{l^{2}}+\frac{J^{2}}{4r^{2}}-2q^{2}\ln{\left(\frac{r}{l}\right)}\equiv-M+\frac{r^{2}}{l^{2}}+\frac{J^{2}}{4r^{2}}-q^{2}\ln{\left(\frac{r^{2}}{l^{2}}\right)}~{},\qquad\Omega(r)=\frac{J}{2r^{2}}~{}.$ (80)
Since we want to write $f(r)$ and $\Omega(r)$ as similar to 2 , we can apply a transformation $\tilde{z}=(r^{2}-r_{+}^{2})/r_{+}^{2}$, which implies
$2q^{2}\ln{\left(\frac{r}{l}\right)}=q^{2}\ln{\left[\frac{r_{+}^{2}(1+\tilde{z})}{l^{2}}\right]}\equiv q^{2}\ln{\left(\frac{r_{+}^{2}}{l^{2}}\right)}+q^{2}\ln{\left(1+\tilde{z}\right)}.$ (81)
In the new coordinate, the event horizon is at $\tilde{z}=0$. For a slowly varying tidal field ($M\omega\ll 1$) and near zone approximation ($M\omega\tilde{z}\ll 1$), we can neglect the terms of order $q^{2}\mathcal{O}(\tilde{z}^{2})$ or $\mathcal{O}(M^{2}\omega^{2}\tilde{z}^{2})$ for a BTZ black hole with small charge ($q\sim\mathcal{O}(M\omega)$). It yields
$2q^{2}\ln{\left(\frac{r}{l}\right)}\sim q^{2}\ln{\left(\frac{r_{+}^{2}}{l^{2}}\right)}+q^{2}\tilde{z}\equiv-q^{2}+q^{2}\ln{\left(\frac{r_{+}^{2}}{l^{2}}\right)}+q^{2}\frac{r^{2}}{r_{+}^{2}}.$ (82)
Now $f(r)$ becomes
$f(r)=-M+\frac{r^{2}}{l^{2}}+\frac{J^{2}}{4r^{2}}-2q^{2}\ln{\left(\frac{r}{l}\right)}\sim-\left[M-q^{2}+q^{2}\ln{\left(\frac{r_{+}^{2}}{l^{2}}\right)}\right]+\left(1-\frac{q^{2}l^{2}}{r_{+}^{2}}\right)\frac{r^{2}}{l^{2}}+\frac{J^{2}}{4r^{2}}.$ (83)
For brevity, we can re-express it as
$f(r)\sim-\tilde{M}+\frac{r^{2}}{\tilde{l}^{2}}+\frac{J^{2}}{4r^{2}}.$ (84)
where the form of $\tilde{M}$ and $\tilde{l}$ are given in 36.